how much does Dr. mann spend on a subway ride each year assuming the current rate of $2.00 per ride and 2 weeks of vacation each year. He makes 16 trips per week

Found 2 solutions by checkley79, jim_thompson5910:
Answer by checkley79(3341)

16(52-2)*2
16*50*2=$1,600 ans.

Answer by jim_thompson5910(35256)

There are 52 weeks in a year. He takes 2 weeks off for vacation. So he works 52-2 = 50 weeks out of the year.

He makes 16 trips per week. So if he does this 50 times, then he makes 16*50 = 800 trips a year.

So he spends a total of 800*2 = 1600 dollars a year.